An alpha particle (q = +2e, m = 4.00 u) travels in a circular path of radius 5.56 cm in a uniform magnetic field with B = 1.76 T. Calculate (a) its speed, (b) its period of revolution, (c) its kinetic energy, and (d)the potential difference through which it would have to be accelerated to achieve this energy.

I just need part c and d
Answer #1

Given :-

  • q = +20    : Charge of alpha particle
  • m = 4.00 u                     : Mass of alpha particle.
  • r = 5.56 cm = 0.0556 m : Radius of circular path
  • B=1.767                       : Magnetic field

Let,

  • \small v    : be the velocity of the alpha particle
  • \small K : be the kinetic energy of the alpha particle
  • c=1.6 x 10-19 C : be the charge
  • \small \Delta V   : be the potential difference.
  • T   : be the period.

Mass m and charge q of the alpha particle in SI unit is,

  • m = 4u = 4 x 1.66 x 10-27 = 6.64 x 10-27 kg
  • q=+2e = 2 x 1.6 x 10-19 = 3.2 x 10-19 J

(a) its speed :-

Velocity of a charged particle, in magnetic field moving in circular path, is given as,

v =\frac{qrB}{m}

v =\frac{3.2 \times 10^{-19} \times 0.0556 \times 1.76}{ 6.64 \times 10^{-27}}

v = 4.72 x 106 m/s

(b) Its period of revolution :-

Period of a charged particle in magnetic field is given as,

T = \frac{2 \pi r}{v}

T = 2 x 3.141 x 0.0556 4.72 x 106

\mathbf{T = 0.74 \times 10 ^{-7} \: s}

(c) Its kinetic energy :-

Kinetic energy is given as,

K =\frac{1}{2} mv^{2}

K =\frac{1}{2}\times 6.64 \times 10^{-27} \times \left ( 4.72 \times10^{6} \right )^{2}

K =3.32 \times 10^{-27} \times 22.28 \times10^{12}

K =73.96 \times 10^{-15} \: J

in eV unit,

\mathbf{K = 4.62 \times 10^{5} \: eV}

This is the kinetic energy of the alpha particle.

(d) The potential difference :-

The potential difference is given as,

\Delta V = \frac{K}{q}

\Delta V = \frac{4.62 \times10^{5} \: \left (eV \right )}{2\left (e \right )}

\mathbf{ \Delta V = 2.31 \times10^{5} \: V }

This is the potential difference through which it would have to be accelerated to achieve this energy.
Answer #1

Given :-

  • q = +20    : Charge of alpha particle
  • m = 4.00 u                     : Mass of alpha particle.
  • r = 5.56 cm = 0.0556 m : Radius of circular path
  • B=1.767                       : Magnetic field

Let,

  • \small v    : be the velocity of the alpha particle
  • \small K : be the kinetic energy of the alpha particle
  • c=1.6 x 10-19 C : be the charge
  • \small \Delta V   : be the potential difference.
  • T   : be the period.

Mass m and charge q of the alpha particle in SI unit is,

  • m = 4u = 4 x 1.66 x 10-27 = 6.64 x 10-27 kg
  • q=+2e = 2 x 1.6 x 10-19 = 3.2 x 10-19 J

(a) its speed :-

Velocity of a charged particle, in magnetic field moving in circular path, is given as,

v =\frac{qrB}{m}

v =\frac{3.2 \times 10^{-19} \times 0.0556 \times 1.76}{ 6.64 \times 10^{-27}}

v = 4.72 x 106 m/s

(b) Its period of revolution :-

Period of a charged particle in magnetic field is given as,

T = \frac{2 \pi r}{v}

T = 2 x 3.141 x 0.0556 4.72 x 106

\mathbf{T = 0.74 \times 10 ^{-7} \: s}

(c) Its kinetic energy :-

Kinetic energy is given as,

K =\frac{1}{2} mv^{2}

K =\frac{1}{2}\times 6.64 \times 10^{-27} \times \left ( 4.72 \times10^{6} \right )^{2}

K =3.32 \times 10^{-27} \times 22.28 \times10^{12}

K =73.96 \times 10^{-15} \: J

in eV unit,

\mathbf{K = 4.62 \times 10^{5} \: eV}

This is the kinetic energy of the alpha particle.

(d) The potential difference :-

The potential difference is given as,

\Delta V = \frac{K}{q}

\Delta V = \frac{4.62 \times10^{5} \: \left (eV \right )}{2\left (e \right )}

\mathbf{ \Delta V = 2.31 \times10^{5} \: V }

This is the potential difference through which it would have to be accelerated to achieve this energy.